OpenStudy (anonymous):

Fernando is going to draw a marble from the bag, replace it, and then draw another marble. What is the probability that Fernando will get a green or blue marble on the first draw and a pink on the second draw? (3 green, 4 yellow, 5 blue, 8 pink)

OpenStudy (ash2326):

@Hpbby total How many marbles are there?

OpenStudy (anonymous):

20 in all

OpenStudy (anonymous):

prob of getting green= 3/20 prob of getting blue =5/20 prob of getting green or blue =(5+3)20=8/20 prob of pink = 8/20..... total prob = 8/20*8/20=4/25